You decide to take advantage of the current online dating craze and start your own Web site. You know that you have 550 people who will sign up immediately and, through a careful marketing research and analysis, determine that membership can grow by 29 percent in the first two years, 21 percent in year 3, and 18 percent in Year 4. How many members do you expect to have at the end of four years?
Answer #1

Ans : Number of members today = 550 members

Number of members that expect at the end of 2 nd year = 550 * ( 1 + growth ) 2

= 550 * ( 1 + 0.29) 2

= 550 * ( 1.29) 2

= 915.255

Numbers of members that expect at the end of 3 nd year = 915.255 * (1 + growth 3 )

= 915.255 * (1+0.21)

= 915.255 * (1.21)

= 1107.45

Number of members that expect at the end of 4 years = 1107.45 * (1+ growth 4)

= 1107.45 ( 1+ 0.18 )

= 1107.45 * 1.18

= 1306.80

Total number of members expected = 1307